OpenStudy (anonymous):

Find the domain of the function: (will input the function when I can use the equation box, lol.) Find the domain of the function: (will input the function when I can use the equation box, lol.) @Mathematics

OpenStudy (anonymous):

\[f(x) = \frac{3+x}{\sqrt{3x^2+2x-5}}\]

OpenStudy (anonymous):

3x^2 +2x -5=0 3x^2+5x-3x-5=0 x(3x+5)-1(3x+5)=0 (x-1)(3x+5) x=1,x=-5/3 Domain=R-{1,-5/3}

OpenStudy (anonymous):

so can it be -5/3? I thought if it's a - under the sqrt symbol it's not possible.

OpenStudy (anonymous):

So domain is x >1 or x < -5/3

OpenStudy (anonymous):

D={x| x > 1 or x<-5/3} or should I state: \[(-\infty,\frac{-5}{3})\cup(1,\infty)\]

OpenStudy (anonymous):

(−∞,−5/3)∪(1,∞) is the domain.